A 27 gallon solution contain 9 gallon acid and 18 gallon water. How much water must be added to get 25% Acid?
Answer by stanbon(75887)

A 27 gallon solution contain 9 gallon acid and 18 gallon water. How much water must be added to get 25% Acid?
--------------
Starter Solution:
amt of solution = 27 gallons ; Amt of acid = 9 gallon
-----------------
Additional solution:
amt = x gallons ; Amt of acid = 0
---------------------
Mixture solution:
Amt = (27+x) gallons ; amt of acid = 0.25(27+x) = 6.75 + 0.25x gallons
-------------------
EQUATION:
acid + acid = acid
9 + 0 = 6.75 + 0.25x
0.25x = 2.25 gallons
x = 9 gallons (# of gallons of water that must be added)
==========================
Cheers,
Stan H.